OpenStudy (anonymous):

how do i find the domain and range for x+5/2x-1. I know the domain should be x is not equal to 1/2. how do i find range?

OpenStudy (anonymous):

have fun

OpenStudy (luigi0210):

set the denominator equal to 0 to find the domain

OpenStudy (anonymous):

but maybe a little in sight would help

OpenStudy (anonymous):

see if you can think of a number right away that this fraction can never be

OpenStudy (anonymous):

luigi: i figured out domain. i don't know how to do range. satellite: confused

OpenStudy (anonymous):

with a little practice you could look at this and say right away that the range cannot include \(\frac{1}{2}\) because...

OpenStudy (anonymous):

the denominator can never be exactly twice the numerator

OpenStudy (anonymous):

it has nothing to do with \(\frac{1}{2}\) being excluded from the domain, that is just a coincidence

OpenStudy (anonymous):

\[\frac{2x+5}{3x-4}\] for example can never be \(\frac{2}{3}\)

OpenStudy (anonymous):

ummm so what are you saying? anything > 1/2 for range?

OpenStudy (anonymous):

no i am saying the range is all numbers except \(\frac{1}{2}\)

OpenStudy (anonymous):

we can do this is a different way if you like, but it is more cumbersome

OpenStudy (anonymous):

ok and the domain all numbers except 1/2 too? i have another problem y=6/x-10, so the domain is everything except 10, would the range be the same 10 for that too?

OpenStudy (anonymous):

no no that is what i was trying to say, that the last one was a coincidence

OpenStudy (anonymous):

you have \[y=\frac{6}{x-10}\] and this has range all real numbers except for 0 why? because a fraction is only zero if the numerator is 0, and your numerator here is 6

OpenStudy (anonymous):

ok so every number greater than 0. so let me try the next y=3x^2-24. Domain is every number except 8. Range is ??? nevermind this is not a fraction so again i'm lost.